# Deploying Rapid Damage Assessments from sUAS Imagery for Disaster Response

### v1

@@ Introduction
There are two known instances of formally documented operational deployments for satellite-based building damage assessment models at disasters. These two instances are the Mississippi Rolling Fork Tornado outbreak (robinson2023rapid), and the 2023 Turkey Earthquake (robinson2023turkey). It is likely that other, undocumented instances of such model deployments exist. In the case of the Mississippi Rolling Fork Tornado outbreak “a U-Net with an ImageNet pre-trained ResNet-50 backbone” was used (robinson2023rapid), and in the case of the 2023 Turkey Earthquake, a “convolutional neural network” was used (robinson2023turkey). While works have proposed alternative architectures, notably with “more than 2000 submissions” being made to the initial xBD competition, there appears to be no other known public documentation indicating operational use of other model architectures (xView2challenge). While it would be convenient to deploy the ML models already developed for satellite imagery in the field for sUAS imagery, previous work has shown that this would be insufficient. First, there would be meaningful degradations in model performance if imagery with a spatial resolution other than that observed during training were passed to such models, which are not currently adapted for the high-resolution imagery generated by sUAS (reed2023scale; manzini2025challenges). Second, if the imagery was downsampled to a spatial resolution that would be usable for such models, the underlying distribution of damage represented in the imagery would be significantly altered, lowering the utility of the collected imagery and potentially misrepresenting the conditions in the scene to decision makers (manzini2025now). While deploying such models on sUAS imagery is possible, a dedicated sUAS solution is necessary.
@@
ResNet101 (he2016deep) DeepLabv3Plus (chen2018encoder) 0.498 0.479
ViT-L ScaleMAE (Pretrained) (reed2023scale) UperNet (xiao2018unified) 0.382 0.365
ResNet101 (he2016deep) PSPNet (zhao2017pyramid) 0.380 0.371
ViT-L ScaleMAE (reed2023scale) UperNet (xiao2018unified) 0.310 0.309
ViT-L ScaleMAE (Pretrained) (reed2023scale) Segmenter (strudel2021segmenter) 0.175 0.164
@@
Attention UNet (oktay2018attention) 0.511 0.492
Vanilla UNet (ronneberger2015u) 0.258 0.249
@@
Following model training, the eight baseline models were evaluated on 12 orthomosaics from four disasters (Hurricanes Michael, Idalia, Musset Bayou Fire, and Mayfield Tornado) containing 4,092 buildings, with the UNet with attention being the most performant with a macro F1 of 0.511. To determine the effectiveness of a model when deployed on an unseen disaster, a disaster-based train/test split, also detailed in (manzini2024crasar), was constructed to evaluate the ability of new models to generalize to “new” disasters. Further, macro F1 was employed to evaluate each model’s predictions across all five damage classes. As the literature notes, predefined building polygons may not align with aerial imagery (manzini2024non; zampieri2018multimodal; vargas2019correcting). To account for this, evaluations were performed on both aligned and unaligned building polygons. Details on alignment in CRASAR-U-DROIDs can be found in (manzini2024non) and (manzini2024crasar). The resultant performances are shown in Table 1. As the Attention UNet model was the most performant, it was selected for deployment.

### v2

@@ Introduction
There are two known instances of formally documented operational deployment for satellite-based building damage assessment models at disasters. These two instances are the Mississippi Rolling Fork Tornado outbreak (robinson2023rapid), and the 2023 Turkey Earthquake (robinson2023turkey). In the Mississippi Rolling Fork Tornado outbreak “a U-Net with an ImageNet pre-trained ResNet-50 backbone” was used (robinson2023rapid), and in the case of the 2023 Turkey Earthquake, a “convolutional neural network” was used (robinson2023turkey). It is likely that other, undocumented instances of such model deployments exist, as similar models are reportedly in use by the United States Defense Innovation Unit (xView2challenge), United Nations Satellite Center (UNOSAT_2025), and Microsoft’s AI for Good Lab (robinson2023rapid; robinson2023turkey). While other alternative architectures exist, notably with “more than 2000 submissions” being made to the initial xBD competition, there appears to be no known public documentation indicating operational use of other model architectures (xView2challenge). While it would be convenient to deploy the ML models already developed for satellite imagery on sUAS imagery, previous work has shown that this would be insufficient. First, there would be meaningful performance degradation if imagery with a spatial resolution other than that observed during training were passed to models not adapted for high-resolution sUAS imagery (reed2023scale; manzini2025challenges). Second, if the imagery was downsampled to a spatial resolution that would be usable for such models, the underlying distribution of damage represented in the imagery would be significantly altered, lowering the utility of the collected imagery and potentially misrepresenting the conditions in the scene to decision makers (manzini2025now). While deploying such models on sUAS imagery is possible, a dedicated sUAS solution is necessary.
@@
ResNet101 (he2016deep) DeepLabv3Plus (chen2018encoder) 0.493 0.475
ViT-L ScaleMAE (Pretrained) (reed2023scale) UperNet (xiao2018unified) 0.372 0.356
ResNet101 (he2016deep) PSPNet (zhao2017pyramid) 0.377 0.368
ViT-L ScaleMAE (reed2023scale) UperNet (xiao2018unified) 0.309 0.303
ViT-L ScaleMAE (Pretrained) (reed2023scale) Segmenter (strudel2021segmenter) 0.174 0.163
@@
Attention UNet (oktay2018attention) 0.554 0.524
Vanilla UNet (ronneberger2015u) 0.531 0.498
@@
Following model training, the eight baseline models were evaluated on 12 orthomosaics from four disasters (Hurricanes Michael, Idalia, Musset Bayou Fire, and Mayfield Tornado) containing 4,092 buildings, with the UNet with attention being the most performant with a macro F1 of 0.554. To determine the effectiveness of a model when deployed on an unseen disaster, a disaster-based train/test split, also detailed in (manzini2024crasar), was constructed to evaluate the ability of new models to generalize to “new” disasters. Further, macro F1 was employed to evaluate each model’s predictions across all five damage classes. As the literature notes, predefined building polygons may not align with aerial imagery (manzini2024non; zampieri2018multimodal; vargas2019correcting). To account for this, evaluations were performed on both aligned and unaligned building polygons. Details on alignment in CRASAR-U-DROIDs can be found in (manzini2024non) and (manzini2024crasar). The resultant performances are shown in Table 1. As the Attention UNet model was the most performant, it was selected for deployment.